The largest measure of an angle of a triangle has a measure that is 10 less than three times the second angle. The measure of the smallest is equal to the difference between the largest and twice the middle angle. Determine the measures of all three angles.
Answer by Benyamin(2)

Let A, B and C be the largest, middle and smallest angles respectively. The assumptions will be modelled as a linear system as follow:
A=3B-10, C=A-2B, A+B+C = 180
Solving this, you can see: A=110, B=40 and C=30.
Good luck